OpenStudy (anonymous):

What are the solutions of 4x2 − 4x − 3 = 0?

OpenStudy (lgbasallote):

you can factor it out..break down the middle term \[4x^2 +2x - 6x - 3 = 0\] group \[(4x^2 + 2x) - (6x + 3) = 0\] factor out \[2x(2x + 1) - 3(2x+1) = 0\] factor again \[(2x+1)(2x-3) = 0\] can you find the solutions now?

OpenStudy (anonymous):

Nice work with the ac method.